OpenStudy (anonymous):

An equation is shown below: 5(2x - 8) + 15 = -15 Write the steps you will use to solve the equation and explain each step.

Mehek (mehek14):

first you would distribute 

Mehek (mehek14):

\(5*2x=10x\\5*(-8)=-40\)

Mehek (mehek14):

so you have \(10x-40+15=-15\)

Mehek (mehek14):

now you would add \(-40+15=-25\)

Mehek (mehek14):

now add \(25+(-15)=10\)

Mehek (mehek14):

so \(10x=10\) divide \(\dfrac{10}{10} = 1\\x=1\)